(4x^3yz^4)(2xy^-2z^-1)(5x^-2y^-2z^3)
simplify

Answer by Edwin McCravy(20056)


(4x3yz4)(2xy-2z-1)(5x-2y-2z3)

Give the y in the 1st factor and the x in the 
2nd factor an exponent of 1:

(4x3y1z4)(2x1y-2z-1)(5x-2y-2z3)

Multiply the coefficients

4򈮁 = 40

Multiply the x's by adding up the exponents

x3x1x-2 = x3+1-2 = x2

Multiply the y's by adding up the exponents

y1y-2y-2 = y1-2-2 = y-3

Multiply the z's by adding up the exponents

z4z-1z3  = z4-1+3 = z6

So you have

40x2y-3z6

The probably want you to use only positive
exponents, so move they y-3 from the
numerator to the denominator as y3

 40x2z6
棗棗棗棗  
   y3

Edwin